# Revealing Nanostructures in High-Entropy Alloys via Machine-Learning Accelerated Scalable Monte Carlo Simulation

### v1

@@ 2.4 Nanostructures in MoNbTaW
Other than FeCoNiAlTi, we also employs our method to sutdy the nanostructure evolution of a well-studied bcc HEA: MoNbTaW. The size of the supercell is $795\times 795\times 795\times 2=1,004,919,750$ atoms. The size of the link-cell is $3\times 3\times 3$, and the energy model is 2nd-NN EPI. The simulation temperatures decrease from 2000 K to 250K, with an temperature inverval of 50 K. For each temperature we run the simulation for 30,000 sweeps. The simulation results is shown in Fig. 5. For T=250K, we can note from the figure is that there are two types of nanostructure: The smaller ones are mainly made up of Nb and Ta, and of a feature size of about 10 nm. It can be seen that the distribution of Nb and Ta elements within relatively random. The shape of these small nanoparticles are also more diverse than that in the FeCoNiAlTi. Other than the small nanostructure, there are also larger nanophases of a feature size of about 100 nm, which is at the same length scale as the supercell. The two nanophases are shown as yellow and purple in the figure. It is easy to see that they are actually the NbW B2 phases. The difference between the yellow and the purple nanophase is that one of them have Ta occupies the center position of the perfect bcc lattice, while the other has Mo occupies the center site. Therefore these two nonophases just correspond to a spontaneous symmetry breaking of an Ising model along either the plus or the minus directions. At T=250 K, a chemical grain-boundary made up of Nb and Ta can be clearly seen between the two nanophases, and the thickness of the chemical grain boundary is about 3 nm. As the temperature increases to 750 K, it can be see that the larger nanophases still exist, which means that those features are more stable against thermal fluctuations. On the other hand, the smaller nanoparticles vanish, along with the chemical grain boundary between the yellow and purple nanophases. The stability of the nanophases presents an interesting phenomenon, which we propose to explain as follows: at low temperatures, the chemical grain boundary acts as a protective barrier, preventing the formation of energetically unfavorable AA or BB nearest-neighbor pairs when the two nanophases come into contact. On the other hand, as the temperature increases, despite the vanish of the chemical grain boundary, the two nanophases also are no longer perfect B2 structures. This disorder reduces the energy cost at the boundary, allowing the nanostructures to persist even in the absence of the protective chemical grain boundary. It would be intriguing if this “disorder-protected phase” observed in our simulations could be directly verified experimentally.
@@ 3 Method
3.1 Implementation of SMC_GPU
For a simple nearest-neighbor 2D Ising model, a widely used parallelization scheme is the checkerboard algorithm. In the checkerboard algorithm, the 2D square lattice is divided into two sublattices, one colored black and the other white. Note that the lattice sites of the same color will not interact with each other, therefore the sites of one color can be updated simultaneously by fixing the sites of the other color. The checkerboard algorithm is very efficient and has been applied to study the 2D and 3D Ising models on different accelerators, including GPU [35, 65, 66], TPU [67], and FPGA [68, 69].
The checkerboard algorithm cannot be directly applied when the Hamiltonian contains interactions beyond the nearest neighbors. Therefore, it is rarely used for studying real materials, where the interaction range is typically beyond nearest neighbors. For such a purpose, a generalization of the checkerboard algorithm can be employed to harness the move-parallelism opportunity. This method is referred to as parallelized link-cell algorithm [70]. In link-cell algorithm, the lattice is decomposed into $N$ cells of fixed size $a_{cell}$, which should be larger than the interaction range. With link-cell algorithm, the serial MC code can be accelerated by up to $N$ times. For a multi-GPU system, the $N$ cells are evenly distributed on all the GPUs, and ghost boundaries are added to take into account the interactions between atoms from different GPUs, as shown in the right part of Fig. 1.
@@ 3.1 Implementation of SMC_GPU
    Parallelization implementation: The actual parallelization implementation on GPUs depends on the energy model. For pairwise model such as EPI, the iteration over the $i_{C}$ index is executed in parallel as CUDA threads on GPU. For more general energy models such as the local_SRO model, the $i_{C}$ index is parallelized in the CUDA blocks dimension, and the calculation of the local energies for each sites within LIZ is parallelized with the CUDA threads, as demonstrated in the flowchart in Fig. 6.
@@
    Link-cell size: Using the link-cell algorithm, the one MC sweep (MC trials over every sites) is decomposed into $n_{A}$ sequential steps, with each one of $n_{C}$ parallel moves. For instance, for a $300\times 300\times 300$ bcc supercell (a total of $000300^{3}\times 2=54,000,$ 000 atoms), assuming next-to-nearest-neighbor interaction, then the link-cell length can be chosen as $3\times a$, where a is the lattice constant. The link-cell size $n_{A}$ is $3\times 3\times 3\times 2=54$, and the number of link-cell is $n_{C}=1,000,000$.
@@
    Search neighbors: Another important consideration is how to efficiently find the neighboring sites, for which their chemical species need to be read from the memory, as shown in the third step in Fig. 6. In practice, we make use of the lattice structure and directly calculate the indices of the neighbors for a specific atom on-the-fly using a list of the relative positions of the neighboring atoms. This method has a constant time complexity, which is advantageous over the more common practice of using k-D tree to store the atomic positions and search the neighbors, which has time complexity of $O(\log(N))$.
Figure 6: A schematic flowchart for the link-cell parallelism algorithm for a machine learning energy model. The outer loop over index $i_{A}$ are executed sequentially, while the inner loop over index $i_{C}$ are executed in parallel as a CUDA kernel since the MC update atoms of the same $i_{A}$ index but different $i_{C}$ indices are independen.
@@ 3.2 The EPI and Local_SRO model
where $\epsilon$ is the uncertainty of the EPI model, $V^{0}$ is the bias term same for all sites, $V_{i}^{p}$ is a single-site term depending only on the chemical component $p$ of atom $i$, $V^{f}$ are the EPI parameters, and $\pi^{f}$ are the number of pair interactions of type $f$. The feature index $f$ is actually made up of three parts $(p,p^{\prime},m)$, representing the element of the local atom, the element of the neighboring atoms, and the coordination shell, respectively. For a canonical system, summing up the local energies over all sites, the total energy is then given by
@@
where $N$ is the total number of atoms and $\Pi_{m}^{pp^{\prime}}$ is the proportion of $pp^{\prime}$ interaction in the $m$-th neighboring shell. Note that due to the fixed chemical composition in canonical system, the single-site term $V_{i}^{p}$ has been absorbed into the constant, and the number of independent EPI parameters is $M(M-1)/2$ for a $M$-component system, which is the reason for the $p^{\prime}<p$ requirement. In practice, the EPI parameters are determined via Bayesian regression [33].
As mentioned, since the EPI model contains only pairwise interactions, the calculation of the energy changes after a MC swap trial is simple: we simply multiply the local energy changes of the two swapping sites by a factor of 2 to take into account the local energy changes of other sites. For more general energy models that contain higher-order/nonlinear interactions, the total energy changes need to be explicitly calculated by adding up the local energies of all sites within the LIZ. As a demonstration of our method for more general machine learning models, here we modify the EPI model by adding a simple quadratic term, and the local energy for each site $E_{i}$ is:
@@
is the percentage of $pp^{\prime}$ interactions for the $m$’s coordination shell, and $n_{m}^{pp^{\prime}}$ is the number of pairs with index $(m,p,p^{\prime})$. Note that we limit the quadratic interaction to the same-element pairs to reduce the number of higher-order terms. The third term on the right of Eq. 6 is a $l^{2}$ regularization of the weights of the quadratic term, reflecting our prior that the higher-order interactions should be a small correction to the EPI model. We name it as the local SRO model since $\Pi_{m}^{pp^{\prime}}$ describes the short-range order within the local chemical environment. It should be noted that unlike EPI, this local SRO model is no longer a linear, or even quadratic function of the pairwise interactions. Instead, it can be expressed as a power series (infinite degree polynomial) of the pairwise interactions due to the $(\pi_{m}^{pp^{\prime}})^{2}$ term. This is also the reason we do not refer to it as “quadratic EPI model”. It is easy to see that the nonlinear interactions can also be modeled with neural networks such as the multilayer perceptrons:
@@
However, due to the large number of principal elements, the number of parameters in such MLP model can easily go beyond the number of data points used in this work (e.g. 220 for FeCoNiAlTi), which lead a high risk of overfitting. Therefore, in this work we will use the local SRO model as a representation of general ML energy models, and leave more complex models for future study when significantly larger DFT datasets are available.
@@ 3.3 DFT dataset
We use the LSMS method [50] to calculate the total energy of the system. LSMS is a all-electron electronic structure calculation method, in which the computational cost scales linearly with respect to the number of atoms. For FeCoNiAlTi, we use a 100 atom supercell with a lattice constant of 6.72 Bohr (0.356 nm). We employs spin-polarized scheme to account for the magnetic interactions in the system. The angular momentum cutoff $l_{max}$ for the electron wavefunctions is chosen as 3, and the default LIZ cutoff radius is chosen as 12.5 Bohr. We used PBE as the exchange-correlational functional. To enhance the representativeness of the DFT data, the total dataset is made up of 120 random generated configurations and 100 configurations from Monte Carlo simulation, as proposed in Ref. [33]. Therefore the total number of configurations is 220, with each one comprising 100 atoms.
For MoNbTaW, the DFT dataset has been reported in Ref. [33], which is also made up of the random samples and the MC samples. The 704 random samples are calculated with supercells of 64, 128, 256, and 512 atoms. The 72 MC samples are obtained from MC simulations at different temperatures, on a 1000-atom supercell. When calculate the energies of the MoNbTaW system, the lattice constant is set at 6.2 Bohr, and the angular momentum cutoff is chosen as 3. The Barth-Hedin local-density approximation are used as the exchange-correlation functional. The local interaction zone (LIZ) is maintained at 59 atoms. To adequately capture the effects of heavier elements in the system, the scalar-relativistic equation is used instead of the conventional Schrödinger equation.
@@ 3.4 MC simulation
Table 3: The MC simulation results obtained with different systems and models. The dimensions are made up of $nx\times ny\times nz\times m$, where $nx,ny,nz$ are the supercell vectors along the $x,y,z$ direction, and $m$ is the number of atoms in the cubic cell.
The results shown in this work are based on five simulation results using different model and systems, as shown in Tab. 3. The MC simulation is a simulated annealing process in the canonical ensemble (NVT). The temperature is initialized as 2000 K and then decrease with a step of 50 K. The initial $\rm{N_{sweeps}}$ are discarded before making measurement or taking the snapshot of the configuration. The number of $\rm{N_{sweeps}}$ generally increases as the temperature decreases to account for the lower acceptance ratio of MC moves. For instance, for $\rm{FeCoNiAlTi\_1M\_EPI}$, the number of sweeps is $4\times 10^{5}$ at 2000 K, then increase to $1.5\times 10^{6}$ at $T=1000K$. In order to extract the nanostructures from the simulation results, we employs the union-find algorithm, which uses a tree structure efficiently manage and manipulate partitions of a set into disjoint subsets. In order to apply the union-find algorithm, we first calculate the local energy of each site, and set a threshold of the energy to determine whether two sites belong to the same set (cluster). Finally, we used VESTA [71] for the visualization of the 3D crystals and clusters.

### v2

@@ 2.4 Nanostructures in MoNbTaW
Other than FeCoNiAlTi, we also employ our method to study the nanostructure evolution of a well-studied bcc HEA: MoNbTaW. The size of the supercell is $795\times 795\times 795\times 2=1,004,919,750$ atoms. The size of the link-cell is $3\times 3\times 3$, and the energy model is 2nd-NN EPI. The simulation temperatures decrease from 2000 K to 250K, with a temperature interval of 50 K. For each temperature, we run the simulation for $10^{5}$ sweeps. The simulation results are shown in Fig. 6. For T=250K, we note that there are two types of nanostructure: The smaller ones are mainly made up of Nb and W, and of a feature size of about 10 nm. It can be seen that the distributions of Nb and W elements are relatively random. The shapes of these small nanoparticles is also more diverse than that in the FeCoNiAlTi. Other than the small nanostructure, there are also larger nanophases of feature sizes of about 100 nm, which are at the same length scale as the supercell. The two nanophases are shown as yellow and purple in the figure. It is easy to see that they are actually the MoTa B2 phases, which is in agreement with previous results [28, 33, 57]. The difference between the yellow and the purple nanophases is that one of them has Ta at the center position of the perfect bcc lattice (denote as B2- in Fig. 6), while the other has Mo at the center site (denote as B2+ in Fig. 6). These two nonophases can be understood as the spontaneous symmetry breaking of an Ising-like model along either the plus or the minus directions. At T=250 K, a chemical grain boundary made up of Nb and Ta can be clearly seen between the two nanophases, and the thickness of the chemical grain boundary is about 3 nm. As the temperature increases to 750 K, it can be seen that the larger nanophases still exist, which means that those features are more stable against thermal fluctuations. On the other hand, the smaller nanoparticles vanish, along with the chemical grain boundary between the yellow and purple nanophases. The stability of the nanophases presents an interesting phenomenon, which we propose to explain as follows: at low temperatures, the disordered A2 boundary acts as a protective barrier, preventing the formation of energetically unfavorable AA or BB nearest-neighbor pairs when the two nanophases come into contact. On the other hand, as the temperature increases, the disordered chemical grain boundary will dissolve into the B2 phases, which introduces additional chemical disorder. This disorder reduces the energy cost at the boundary, allowing the nanostructures to persist even in the absence of the protective chemical grain boundary. It would be intriguing if this “disorder protected nanophase” revealed in the simulations could be directly verified by future experiments.
@@ 3 Method
3.1 Theoretical speedup of SMC vs traditional methods
The theoretical speedup of the SMC method as compared to other methods are estimated as follows: For DFT and linear-DFT, we assume that a 100-atoms SCF calculation takes one hour, based on our experience with the MuST-KKR code [66]. The speed of the GNN method (Allegro) is estimated from the $\rm{Li_{3}PO_{4}}$ structure of 421,824 atoms in Ref. [24]. Note that an MC sweep is defined as making an MC trial move over each lattice site. Combining it with the intrinsic $O(N^{3})$ scaling in the DFT method, the total computational cost of DFT then scales as $O(N^{4})$, as signified in Fig. 1 (d). Building on the principles of near-sightedness [67], the LSMS method employs an approximation that confines electron scattering to a local interaction zone. This approximation reduces the computational cost of energy evaluation to linear scaling, therefore the total computational cost for an MC sweep scales approximately as $O(N^{2})$. For a one million-atom system, this reduces the computational cost by a factor of $10^{8}$, as shown in the orange line in Fig. 1 (d). Despite the enhanced scaling behavior, linear-DFT still requires solving the Schrödinger equations explicitly, which contributes a large prefactor to the total computational cost. By replacing the computationally expensive DFT method with machine learning models such as the GNN model, the computational cost can be further reduced by a factor of $10^{8}$, as shown in the green line in Fig. 1 (d). The speedup from the SMC-GPU implementation can be divided into multiple parts: First, the SMC parallelization strategy replaces the energy evaluation of the whole system with calculating the energy changes of the LIZ, which is independent of the system size. This strategy reduces the computational complexity from $O(N^{2})$ to $O(N)$, as illustrated in Fig. 1 (d). Second, the SMC method can simultaneously harness the thousands or tens of thousands of cores in a modern high-performance GPU (e.g. 14,592 FP32 CUDA cores in an NVIDIA H800 GPU). Third, the energy models used in this work are relatively simple compared to the GNN models. Based on the preceding discussion, the total acceleration ratio of the SMC-GPU method relative to a CPU-based DFT method, for a system of $N$ atoms, can be decomposed into contributions from nearsightedness $S_{NS}$, ML acceleration $S_{ML}$, and SMC implementation $S_{SMC-GPU}$, and written as:
$S=S_{NS}\times S_{ML}\times S_{SMC-GPU}\approx 10^{7}\times N^{3},$ (1)
as illustrated in Fig. 1 (d).
3.2 Implementation of SMC-GPU
For a simple nearest-neighbor 2D Ising model, a widely used parallelization scheme is the checkerboard algorithm. In the checkerboard algorithm, the 2D square lattice is divided into two sublattices, one colored black, and the other white. Note that the lattice sites of the same color will not interact with each other; therefore, the sites of one color can be updated simultaneously by fixing the sites of the other color. The checkerboard algorithm is very efficient and has been applied to study the 2D and 3D Ising models on different accelerators, including GPU [35, 68, 69], TPU [70], and FPGA [71, 72].
The checkerboard algorithm cannot be directly applied when the Hamiltonian contains interactions beyond the nearest neighbors. Therefore, it is rarely used for studying real materials, where the interaction range is typically beyond nearest neighbors. For such a purpose, a generalization of the checkerboard algorithm can be employed to harness the move-parallelism opportunity. This method is referred to as a parallelized link-cell algorithm [73]. In the link-cell algorithm, the lattice is decomposed into $n_{C}$ cells of fixed size $a_{cell}$, which should be larger than the interaction range. With the link-cell algorithm, the serial MC code can be accelerated up to $n_{C}$ times. For a multi-GPU system, the $n_{C}$ cells are evenly distributed on all the GPUs, and ghost boundaries are added to take into account the interactions between atoms from different GPUs, as shown in Fig. 1 (b).
@@ 3.2 Implementation of SMC-GPU
    Parallelization implementation: The actual parallelization implementation on GPUs depends on the energy model. For pairwise models such as EPI, the iteration over the $i_{C}$ index is executed in parallel as CUDA threads on the GPU. For more general energy models such as the local_SRO model, the $i_{C}$ index is parallelized in the CUDA blocks dimension, and the calculation of the local energies for each site within LIZ is parallelized with the CUDA threads, as demonstrated in the flowchart in Fig. 7.
@@
    Link-cell size: Using the link-cell algorithm, the one MC sweep (MC trials over every sites) is decomposed into $n_{A}$ sequential steps, with each one of $n_{C}$ parallel moves. For instance, for a $300\times 300\times 300$ bcc supercell (a total of $000300^{3}\times 2=54,000,$ 000 atoms), assuming next-to-nearest-neighbor EPI interaction, then the link-cell length can be chosen as $3\times a$, where a is the lattice constant. The link-cell size $n_{A}$ is $3\times 3\times 3\times 2=54$, and the number of link-cell is $n_{C}=1,000,000$.
@@
    Search neighbors: Another important consideration is how to efficiently find the neighboring sites, for which their chemical species need to be read from the memory, as shown in the third step in Fig. 7. In practice, we make use of the lattice structure and directly calculate the indices of the neighbors, as well as the features of the local chemical environment, on-the-fly using a list of the relative positions of the neighboring atoms. This method has a constant time complexity, which is advantageous over the more common practice of using k-D tree to store the atomic positions and search the neighbors, which has a time complexity of $O(\log(N))$.
Figure 7: A schematic flowchart for the link-cell parallelism algorithm for a machine learning energy model. The outer loop over index $i_{A}$ is executed sequentially, while the inner loop over index $i_{C}$ is executed in parallel as a CUDA kernel since the MC update atoms of the same $i_{A}$ index but different $i_{C}$ indices are independent.
@@ 3.3 The EPI and Local_SRO model
where $\epsilon$ is the uncertainty of the EPI model, $V^{0}$ is the bias term same for all sites, $V_{i}^{p}$ is a single-site term depending only on the chemical component $p$ of atom $i$, $V^{f}$ are the EPI parameters, and $\pi^{f}$ are the number of pair interactions of type $f$. The feature index $f$ is actually made up of three parts $(p,p^{\prime},m)$, representing the element of the local atom, the element of the neighboring atoms, and the coordination shell, respectively. For a system of fixed chemical concentrations, summing up the local energies over all sites, the total energy is then given by
@@
where $N$ is the total number of atoms and $\Pi_{m}^{pp^{\prime}}$ is the proportion of $pp^{\prime}$ interaction in the $m$-th neighboring shell. Note that due to the fixed chemical concentration in the canonical system, the single-site term $V_{i}^{p}$ has been absorbed into the constant, and the number of independent EPI parameters is $M(M-1)/2$ for a $M$-component system, which is the reason for the $p^{\prime}<p$ requirement. In practice, the EPI parameters are determined via Bayesian regression [33] or stochastic gradient descent.
As mentioned, since the EPI model contains only pairwise interactions, the calculation of the energy changes after an MC swap trial is simple: we simply multiply the local energy changes of the two swapping sites by a factor of 2 to take into account the local energy changes of other sites. For more general energy models that contain higher-order/nonlinear interactions, the total energy changes need to be explicitly calculated by adding up the local energies of all sites within the LIZ. As a demonstration of our method for more general machine learning models, here we modify the EPI model by adding a simple quadratic term, and the local energy for each site $E_{i}$ is:
@@
is the percentage of $pp^{\prime}$ interactions for the $m$’s coordination shell, and $n_{m}^{pp^{\prime}}$ is the number of pairs with index $(m,p,p^{\prime})$. Note that we limit the quadratic interaction to the same-element pairs to reduce the number of higher-order terms. The third term on the right of Eq. 6 is a $l^{2}$ regularization of the weights of the quadratic term, reflecting our prior that the higher-order interactions should be a small correction to the EPI model. We name it the local_SRO model since $\Pi_{m}^{pp^{\prime}}$ describes the short-range order within the local chemical environment. It should be noted that, unlike EPI, this local SRO model is no longer a linear or even quadratic function of the pairwise interactions. Instead, it can be expressed as a power series (infinite degree polynomial) of the pairwise interactions due to the $(\pi_{m}^{pp^{\prime}})^{2}$ term. This is also the reason we do not refer to it as a “quadratic EPI model”. It is easy to see that the nonlinear interactions can also be modeled with neural networks such as the multilayer perceptrons:
@@
However, due to a large number of principal elements, the number of parameters in such an MLP model can easily go beyond the number of data points used in this work (e.g. 220 for FeCoNiAlTi), which leads to a high risk of overfitting. Therefore, in this work we will use the local SRO model as a representation of general ML energy models, and leave more complex models for future study when significantly larger DFT datasets are available.
@@ 3.4 DFT dataset
We use the LSMS method [50] to calculate the total energy of the system. LSMS is an all-electron electronic structure calculation method, in which the computational cost scales linearly with respect to the number of atoms. For FeCoNiAlTi, we use a 100-atom supercell with a lattice constant of 6.72 Bohr (0.356 nm). We employs a spin-polarized scheme to account for the magnetic interactions in the system. The angular momentum cutoff $l_{max}$ for the electron wavefunctions is chosen as 3, and the LIZ is chosen as 86 atoms. We used PBE as the exchange-correlational functional. To enhance the representativeness of the DFT data, the total dataset is made up of 120 random generated configurations and 100 configurations from Monte Carlo simulation, as proposed in Ref. [33]. Therefore the total number of configurations is 220, with each one comprising 100 atoms.
For MoNbTaW, the DFT dataset has been reported in Ref. [33], which is also made up of the random samples and the MC samples. The 704 random samples are calculated with supercells of 64, 128, 256, and 512 atoms. The 72 MC samples are obtained from MC simulations at different temperatures, on a 1000-atom supercell. When calculate the energies of the MoNbTaW system, the lattice constant is set at 6.2 Bohr, and the angular momentum cutoff is chosen as 3. The Barth-Hedin local-density approximation are used as the exchange-correlation functional. The local interaction zone is maintained at 59 atoms. To adequately capture the effects of heavier elements in the system, the scalar-relativistic equation is used instead of the conventional Schrödinger equation.
@@ 3.5 MC simulation
Table 3: The systems and models used for the MC simulation. The dimensions are made up of $nx\times ny\times nz\times m$, where $nx,ny,nz$ are the supercell vectors along the $x,y,z$ direction, and $m$ is the number of atoms in the cubic cell. The times of one MC sweep $\rm{T_{sweep}}$ for different cases are also shown to evaluate the speed.
The results shown in this work are based on five simulation results using different models and systems, as shown in Tab. 3. The MC simulation is a simulated annealing process in the canonical ensemble (NVT). The temperature is initialized as 2000 K and then decrease with a step of 50 K. The initial $\rm{N_{sweeps}}$ sweeps are discarded before making measurement or taking a snapshot of the configuration. The number of $\rm{N_{sweeps}}$ generally increases as the temperature decreases to account for the lower acceptance ratio of MC moves. For instance, for $\rm{FeCoNiAlTi\_1M\_EPI}$, the number of sweeps is $4\times 10^{5}$ at 2000 K, then increase to $1.5\times 10^{6}$ at $T=1000K$. For $\rm{FeCoNiAlTi\_1B\_EPI}$, we evenly decrease the simulation temperature from 2000 K to 1000 K using a step of 50 K. For the 1B atom system, we no longer record the MC configurations and run $10^{5}$ MC sweeps before recording the configuration at $T=1000$ K. Note that, generally speaking, the $10^{5}$ number of MC sweeps is not large enough for a supercell as large as one billion atoms to reach thermal equilibrium, so the simulation is closer to simulated annealing. Nevertheless, this nonequilibrium does not necessarily present as a problem since the actual synthesis of alloys is typically a nonequilibrium process involving a rapid heating-cooling process, in which the phases at high-temperature can be trapped by the kinetic barrier and be maintained at low temperatures.
In order to extract the nanostructures from the simulation results, we employs the union-find algorithm, which uses a tree structure to efficiently manage and manipulate partitions of a set into disjoint subsets. In order to apply the union-find algorithm, we first calculate the local energy of each site and set a threshold of the energy to determine whether two sites belong to the same set (cluster). MC simulations are performed on a workstation with two 80 GB NVIDIA H800 SXM5 GPUs that are connected via 400 GB/s bidirectional bandwidth NVLink. In practice, we find that one GPU is large enough to fit the one-billion atom system, and the two GPUs are simultaneously used via either temperature parallelization or lattice decomposition, as described in section. 2.1. For visualization, we used VESTA [74] to generate the 3D crystals and nanoparticles figures.